The top of a 4.5m long inclined plane is 2.6m above the horizontal. An 8.0kg load slides down the inclined, its motion impeded by a 21.1N frictional force.
Calculate (i) the acceleration of the load down the plane
(ii) the coefficient of kinetic friction between the load and the plane.
1
Expert's answer
2020-07-06T15:40:59-0400

Calculate the angle that the incline makes above the horizontal:

"\theta=\text{arcsin}(h/l)=35°."

According to Newton's second law, if we point the x-axis down the incline, the acceleration can be expressed as


"ma=mg\text{ sin}\theta-f,\\
a=g\text{ sin}\theta-f/m=3\text{ m/s}^2."

The frictional force is the normal force times coefficient of friction:


"f=\mu N,\\
N=mg\text{ sin}\theta,\\
\mu=f/(mg\text{ sin}\theta)=0.47"